OpenStudy (anonymous):

I'm am trying to find the thickness of aluminum foil. the length, width, and weight of it was given. What steps do I need to take to find the thickness?

OpenStudy (anonymous):

First you need to know what the mass density of aluminum is. We will just call it rho for now. You know its mass m = w/g and you know the area, A = length times width, of your sample. If we call x the thickness of the sample, then its volume is Ax, so its mass m = rho*Ax. Solve this equation for x: x = m/(rho*A),

OpenStudy (anonymous):

So the mass is 5g; I'm going to multiply the mass by the area to get the thickness?

OpenStudy (anonymous):

What you do is multiply the area of the sample by its density, then divide that product into the mass to calculate the thickness.

OpenStudy (anonymous):

Its thickness should be something in the neighborhood of 3 to 6 thousandths of an inch to give you a ballpark figure.